Solve the system using Gaussian elimination with back-substitution.

3x1 + 2x2 -x3= -15
5x1+ 3x2 + 2x3 =0
3x1+ x2 +3x3=11
11x1+7x2 = -30
1
Expert's answer
2020-12-21T15:28:15-0500

3x1 + 2x2 -x3= -15

5x1+ 3x2 + 2x3 =0

3x1+ x2 +3x3=11

11x1+7x2 = -30

"\begin {bmatrix}
 3 & 2 & -1 \ \ \ \ \ \ -15 \\
 5 & 3 & 2 \ \ \ \ \ \ \ \ \ \ \ \ \ 0\\
 3 & 1 & 3 \ \ \ \ \ \ \ \ \ \ \ \ 11 \\
 11 & 7 & 0 \ \ \ \ \ \ \ -30
\end{bmatrix}
=" "\begin {Vmatrix}
R_2 =5R_1-3R_2 \\
R_3 = R_1-R_3 \\
R_4 = 5R_4-4R_2
\end{Vmatrix} 
=" "\begin {bmatrix}
 3 & 2 & -1 \ \ \ \ \ \ -15 \\
 0 & 1 & -1 \ \ \ \ \ \ \ -75\\
 0 & 1 & -2 \ \ \ \ \ \ \ \ \ \ \ 26 \\
 0 & 2 & -22 \ \ \ \ \ \ \ -150
\end{bmatrix}="


"\begin {Vmatrix}
R_1 =R_1-2R_2 \\
R_2 =R_2-R_3 \\
R_3 = 2R_3-R_4 \\
R_4 = 2R_2- R_4
\end{Vmatrix} 
=" "\begin {bmatrix}
 3 & 0 & 11 \ \ \ \ \ \ -135 \\
 0 & 1 & -1 \ \ \ \ \ \ \ -75\\
 0 & 0 & 18 \ \ \ \ \ \ \ \ \ \ \ 202 \\
 0 & 0 & 20 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 0
\end{bmatrix}=" "\begin {Vmatrix}
R_1 =\frac{R_1}{3} \\
R_2 =R_2-R_3 \\
R_3 = \frac{R_3}{18} \\
R_4 = 10R_3-9R_4
\end{Vmatrix} 
="


"\begin {bmatrix}
 1 & 0 & \frac{11}{3} \ \ \ \ \ \ -\frac{135}{3} \\
 0 & 1 & -1 \ \ \ \ \ \ -75\\
 0 & 0 & 1 \ \ \ \ \ \ \ \ \ \ \frac{101}{9} \\
 0 & 0 & 0 \ \ \ \ \ \ \ \ \ 2020
\end{bmatrix};"


In "R_4" "0 \ne 2020" . So this system has no solution.